David and Kate can put aside $1,500 as savings each month. They have managed to save $ 50,000 so far during their working lives and this money is earning 3.15% annual interest, compounded monthly, in a bank fixed deposit. As self - employed they have no KiwiSaver funds. David and Kate are renting at present but wish to purchase a home. Their goal is to put down a 20% deposit on a $ 700,000 home in three years’ time. They are finding, however, that average 10% p/a house prices rise, is making it difficult as they are chasing a moving target.

The couple also has a credit card on which they owe $6,200 according to their last bill. They are trying not to use the card anymore but tend to only pay off the minimum amount (2% of the balance) each month. The interest rate on the card is 21.74%, compounded monthly. In addition, the couple also owe $7,000 to their bank from taking a small business loan to fund their purchase of the equipment in their shop, they are paying an interest rate of 6.5% for the loan at the moment.

His friend has advised him that his BitCoin investments have returned 45% p/a over the last two years. David hopes he will be able to persuade Kate to invest their $ 50,000 in Bitcoins for a period of two years, while investing any additional future monthly savings in the bank.

Jeremy recommends that they invest in an IT start - up established by a friend of his, a just listed firm called SMARTIT . SMARTIT shares cost $5.15 each and the prospectus promises a dividend of 15c per quarter, as well as forecasting capital growth of 20% per year. Being a start - up, SMARTIT has never made a profit in its 3 years of existence. Its shares have changed in price by 22%, - 15%, and 35% over the last three years.

Assuming the rates of return indicated in the case study could be achieved, and did not change over the next three years, would any of the three alternative investment plans enable them to achieve their goal of saving a deposit for a home?
Answer #1

Savings per month = $1500 per month

Total Savings in 3 years i.e 36 months = $54000

Investment in FD = $ 50000

Value with interst @ 3.15% monthly compounding is given by

A = P(1+r/n)^nt

So, A = 50000*(1+0.0315/12)^(12*3)

A = 54948.65

Deposit for House = 700000*20%

= 140000

Amount required after 3 years @ 10% increase every year = 140000 *(1+.1)^3

= 186340

Credit Card Liability

Month Liability Interest Payment Balance Month Liability Interest Payment Balance 126 6186 126 6172 126 6158 125 6144 125 6131 125 6117 125 6103 124 6089 1246076 5954 5941 5927 5914 5901 108 108 107 107 107 107 106 106 106 106 5941 5927 5914 5901 2 6186 3 6172 6158 5 6144 6131 7 6117 6103 9 6089 6076 2 5874 5874 5861 2 5861 5835 12 6049 136035 16021 123 6049 1236035 123 6021 5796 5783 5770 5757 5796 5783 5770 5757 105 105 16 5994 175981 18 5968 3 122 5994 1225981 122 5968 1225954 5731 5731 5718 Total 2230 1905 2141 Total Payment 4371

Loan Amount = 7000

Interest payable for 3 years = 7000*6.5%*3 years = $ 1365

Option 1

Total Savings = 54000+54948.65 = 108948.65

Total Cost = 4370.73+1365 = 5735.73

Net Savings = 103212.93

Option 2

Investment in Bitcoin = 50000

Return in 2 Years = 45000

Savings in Bank = 54000

Interest on Savings = 95000*(1+0.0315/12)^12 = 98036.08

Total Savings = 98036.08+54000 = 152036.08

Total Cost = 4370.73+1365 = 5735.73

Net Savings = 146300.36

Option 3

Dividend = 0.15*12 qtrs = 1.8

No. of Shares = 50000/5.15

= 9708 shares

Total Dividend =17474.4

Value of share after 3 years = 50000*(1+0.2)^3 = 86400

Total Savings = 140400

Total Cost = 5737.53

Net Savings = 134664.27

So, any of the option will not be able to save the down payment for the house.